OpenStudy (anonymous):

how do you graph this equation: y + 3 = -3(x + 2)

OpenStudy (anonymous):

sorry, its actually y + 4 = -3(x + 2)

OpenStudy (anonymous):

It's in point-slope form (fortunately).\(y-y_0=m(x-x_0)\) means you have a line with slope \(m\) that intersects the point \((x_0,y_0)\).